Express the following angles in degrees, revolution and radians

(a) 45o

(b) 65 rad

(c) 120 rev


1
Expert's answer
2021-05-23T16:34:14-0400

(a)

"45^{\circ}\cdot\dfrac{1\ rev}{360^{\circ}}=0.125\ rev,""45^{\circ}\cdot\dfrac{\pi\ rad}{180^{\circ}}=0.78\ rad."

(b)

"65\ rad\cdot\dfrac{180^{\circ}}{\pi\ rad}=3724^{\circ},""65\ rad\cdot\dfrac{1\ rev}{2\pi\ rad}=10\ rev."

(c)

"120\ rev\cdot\dfrac{360^{\circ}}{1\ rev}=43200^{\circ},""120\ rev\cdot\dfrac{2\pi\ rad}{1\ rev}=754\ rad."